OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 5x + 4 = 0 Part A: Describe the solution(s) to the equation by just determining radicand. show your work. Part B: Solve 4x2 -12x + 5 = 0 using an appropriate method. Show the steps of your work, and explain why you chose the method used. Part C: Solve 2x2 -10x + 3 = 0 by using a method different from the one you used in Part B. Show the steps of your work. Someone help please, Thanks!

OpenStudy (anonymous):

@mathstudent55 @sleepyjess @ganeshie8 @pooja195 @ParthKohli

Nnesha (nnesha):

what is the radicand ? do you know ?

OpenStudy (anonymous):

No, I have no idea

Nnesha (nnesha):

do you know the quadratic formula ?

OpenStudy (anonymous):

Uhmmm @Nnesha I don't think so

OpenStudy (anonymous):

Sorry, I was getting a drink of water lol

Nnesha (nnesha):

then i think you haven't take the notes and directly jumped to the questions anyways the quadratic formula is \[\large\rm x=\frac{-b \pm \sqrt{\color{Red}{b^2-4ac}}}{2a} \] the value under the root known as radicand b^2-4ac radicand also known as discriminant

Nnesha (nnesha):

\(\huge\color{reD}{\rm b^2-4ac}\) `Discriminant` you can use this to find if the equation is factorable or not/or to find numbers of solutions if ` b^2-4ac > 0` then there are 2 real zeros if ` b^2-4ac = 0` then there is one real root if ` b^2-4ac < 0` then you will get two complex roots (no -x-intercept)

OpenStudy (anonymous):

Wait, so how do I answer Part A @Nnesha ?

Nnesha (nnesha):

first find the the radicand value of the given equation by plugging in b ,c values and then read what i said n my last comment to *describe * the solutions

Nnesha (nnesha):

\[\large\rm Ax^2+Bx+C=0\] quadratic equation where a= leading coefficient b= coefficient of x term c= constant term

Nnesha (nnesha):

x^2 + 5x + 4 = 0 what's a b and c ??

OpenStudy (anonymous):

@Nnesha a is 0, b is 5, and c is 4

OpenStudy (anonymous):

b^2 - 4ac = 5^2 - 4(0)(4)

OpenStudy (anonymous):

= 25, correct?

Nnesha (nnesha):

if a is 0 then 0x^2 should be the same as 0 times x^2=0

Nnesha (nnesha):

it's 1x^2+5x+4=0 a =1 not 0

OpenStudy (anonymous):

Oh right, a is 1. I forgot that x^2 is the same as 1x^2

Nnesha (nnesha):

1 is invisible not 0

OpenStudy (anonymous):

Right

Nnesha (nnesha):

correct substitute again

OpenStudy (anonymous):

5^2 - 4(1)(4) = 25 - 16 = 9, correct?

Nnesha (nnesha):

correct now how would you describe that ?

OpenStudy (anonymous):

What do you mean by describe?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \(\huge\color{reD}{\rm b^2-4ac}\) `Discriminant` you can use this to find if the equation is factorable or not/or to find numbers of solutions if ` b^2-4ac > 0` then there are 2 real zeros if ` b^2-4ac = 0` then there is one real root if ` b^2-4ac < 0` then you will get two complex roots (no -x-intercept) \(\color{blue}{\text{End of Quote}}\)

Nnesha (nnesha):

i mean how would you describe solution if the radicand is 9 then what type of the solution you will get ?

OpenStudy (anonymous):

You will get 2 zeroes?

Nnesha (nnesha):

correct 2 *real* roots

OpenStudy (anonymous):

one is x and one is y?

Nnesha (nnesha):

now the solutions( x-intercepts) where line intersect the x-axis when y=0